The stock of Payout Corp. will go ex-dividend tomorrow. The dividend will be $0.80 per share, and there are 22,000 shares of stock outstanding. The market-value balance sheet for Payout is shown below. Ignore taxes.

Assets   Liabilities and Equity

Cash    $200,000   Equity $1,210,000

Fixed assets $1,010,000

a. What price is Payout stock selling for today? (Round your answer to 2 decimal places.)

b. What price will it sell for tomorrow? (Round your answer to 2 decimal places.)
Answer #1

a

Price = MV of equity/shares = 1210000/22000 = 55

b

price tomorrow = price today-dividend = 55-0.8 = 54.2